OpenStudy (anonymous):

evaluate trig expression sin(81pi/4) tan(-11pi/3) and sec(17pi/6) how do I evaluate these expressions?

OpenStudy (anonymous):

sin (81pi/4) = sin (pi/4) = \(\dfrac{\sqrt{2}}{2}\) tan (-11pi/3) = tan (pi/3) = \(\sqrt{3}\) sec (17pi/6) = sec (5pi/6) = \(\dfrac{-2\sqrt{3}}{3}\) replace all to get the answer

OpenStudy (joannablackwelder):

Do you know how to find coterminal angles?

OpenStudy (joannablackwelder):

And do you know the special right triangles? 30-60-90 and 45-45-90?

OpenStudy (anonymous):

no how do you find the angles?

OpenStudy (anonymous):

also i thought that the tan one was 2pi/3 why is it pi/3?

OpenStudy (anonymous):

I give you the first one, Ok 81pi/4 from pi/4 

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

now, I just count